Time Value of Money Concept

The following situations involve the application of the time value of money concept. Use the full factor when calculating your results.

Use the appropriate present or future value table:

FV of $1, PV of $1, FV of Annuity of $1 and PV of Annuity of $1

1. Janelle Carter deposited $9,540 in the bank on January 1, 2000, at an interest rate of 10% compounded annually. How much has accumulated in the account by January 1, 2017? Round to the nearest whole dollar.

2. Lee Spony made a deposit in the bank on January 1, 2010. The bank pays interest at the rate of 11% compounded annually. On January 1, 2017, the deposit has accumulated to $17,930. How much money did Lee originally deposit on January1, 2010? Round to the nearest whole dollar.
Answer #1

Part 1

FV = PV *(1+r)^n = 9540*(1+10%)^17 = $48220 (using formula)

FV = PV * Future value of $1 n= 17, i = 10% = 9540*5.05447= $48220 (using table value)

Part 2

PV = future value / (1+r)^n = 17930/(1+11%)^7 = $8636 (using formula)

PV = future value / future value of $1 n=7 i=11% = 17930/2.07616 = $8636 (using table value)